# A Photon Cloud Induced from an Axion Cloud

### v1

@@ Abstract
It is known that the axion-photon coupling can lead to quantum stimulated emission of photons and classic exponential amplification of electromagnetic (EM) fields at half the axion mass frequency, when the axion density or the coupling constant is sufficiently large. In this work, we studied the EM photon cloud induced from an axion cloud around a Kerr black hole in the first order of the coupling constant classically. In the presence of a static EM background (like Wald extended solution) valid in realistic astrophysical environment, we found that an EM photon cloud emerges, oscillating at the same frequency as the axion cloud and growing exponentially in accordance with the axion cloud when the superradiant condition for the axion field is satisfied. Notably, the first-order study indicates that this photon cloud exists with no threshold for the coupling constant. The evolution of the EM photon cloud with time and azimuthal angle is obtained analytically while the cross-sectional distribution is solved numerically. Intriguingly, the induced EM field exhibits significantly different symmetries in contrast to the background EM field, which may serve as an indication of the existence of both an axion cloud and axion-photon coupling. Furthermore, in the near horizon region, the induced EM field has some overlap with the photon region as expected.
@@ I Introduction
Recently there are extensive experimental and observational efforts to find axionic imprints. Superradiant mechanisms can give us information about their existance as axions form dense clouds around spinning black holes [9]. The superrradiance mechanism is a very well known method of extracting energy from a rotating black hole [10, 11, 12]. It was shown by Penrose [13] that if a particle accretes into a Kerr black hole then energy can be extracted from the black hole. It was shown by Misner [14] that there is a constrain of the frequency of scatter waves off the black hole to be less than the rotational frequency of the black hole. Then if this constrained condition is satisfied it was shown by Teukolsky [15] that this energy extraction mechanism also works in the case of EM and gravitational waves, though the Kerr solution remains stable under massless scalar, EM and gravitational perturbations without reflection.
It is well known that an isolated black hole cannot have an EM field unless it is endowed with a net electric charge. Therefore, for an isolated black hole to have some EM effects there should be a mechanism for charging it. If a black hole is in an astrophysical environment with a galactic magnetic field, EMs produced by external sources may charge it. In [16] an axisymmetric black hole which was placed in an originally uniform magnetic field and no significant field enhancement effects were found. However, if the black hole is rotating, the rotational effects of the black hole, produce electric fields near the black hole. Extraction of energy and charge from a black hole was studied in [17]. The superradiance effects and the multimessenger signals of an EM field were studied in [18, 19].
It was shown in [20] that the presence of light scalars like axions can trigger superradiant instabilities around Kerr BHs and lead to the deposition of the black hole rotation energy in a cloud. Also it was shown in [21] that the presence of an axion field can lead to the reversal of rotation in Kerr black holes. It was also shown that instabilities are generated and a fraction of the cloud energy is transferred to EM blasts. The resulting large energy density of the cloud rotates around the black hole, and results to a monochromatic gravitational wave radiation that drain the cloud over parametrically longer times [22].
Although the axion-photon coupling is expected to be very weak, it was suggested that the quantum stimulated emission of photons can be generated at large enough axion number [23, 24], and conjectured that blasts of light could be emitted from black hole systems [25, 26]. Recently at classical level, it was shown that EM fields can be exponentially amplified in Kerr-axion cloud system, when the amplitude of the axion cloud multiplied by the coupling constant is larger than a critical value [20]. Even in the presence of plasma, the EM instability is still controlled by the axionic coupling [9]. In [27], the origin of the instability from the axion-photon coupling was analyzed in flat space through the Mathieu equation, and it was found that a background EM field can produce an axion profile around a Kerr black hole perturbatively. On the other side, from the European Horizon Telescope (EHT) observation of the polarimetric measurements of M87*, stringent constraints on the axion-photon coupling constant were given in the mass range of $\sim\left(10^{-21}-10^{-20}\right)$eV from the azimuthal distribution of electric vector position angle [28]. Furthermore, the deviation of photon geodesics due to oscillating metric perturbations induced from the superradiant clouds was predicted in [11].
@@
In this work we studied the EM photon cloud induced from an axion cloud via the axion-photon coupling surrounding a Kerr black hole. In the first order of the coupling constant, we found that a background EM field is necessary to have a first-order induced EM structure, and if it is stationary then the frequency of the induced photon cloud will be the same as the frequency of the axion cloud. It means that it grows or decays in accordance with the axion cloud when the superradiant condition for the axion field is satisfied or not, while the oscillating frequency is also the same. It is well known that the timescale of the superradiant instability is much larger than the timescale of the oscillation, therefore we can ignore the growth of the amplitude and focus on the oscillating part. It is surprising that even with very small couplings, no threshold is needed, there exists an oscillating photon cloud at the same frequency as the axion cloud and it will grow exponentially when the axion cloud is growing. Whereas in the previous studies, without a background EM field, a threshold for the amplitude of the axion cloud multiplied by the coupling constant and an initial EM pulse are both required, otherwise there is no instability and the initial EM fluctuation will decay exponentially. In contrast, for the dark photon-photon coupling case, a dark photon cloud can transfer its energy to form a rotating EM field directly [18].
@@
The work is organized as follows. Firstly in Sec. II we study the basic field equations in the first-order of the coupling constant, then in Sec. III we show how an EM photon cloud is induced from the axion cloud in the presence of an EM background and discuss symmetries. Moreover, in Sec. IV we study the induced electric and magnetic fields respectively, and in Sec. V we analyze the symmetries of most quantities and confirm the parity violation. Finally in Sec. VI we conclude.
@@ II First-order Analysis of the Axion-Photon Coupling
$\mathcal{L}_{m}=-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}-\frac{1}{2}\nabla_{\mu}\phi\nabla^{\mu}\phi-\frac{1}{2}\mu_{1}^{2}\phi^{2}-\frac{k_{a}}{2}\phi F_{\mu\nu}{}^{*}F^{\mu\nu}~{},$ (2)
[…]
@@
$\nabla^{\mu}\nabla_{\mu}\phi-\mu_{1}^{2}\phi-\frac{k_{a}}{2}F_{\mu\nu}{}^{*}F^{\mu\nu}=0~{},$ (3)
@@
In our study, the Kerr geometry is fixed as the background, since the backreactions of both axion cloud and the EMs on the metric are in the second-order of the fluctuations. We expand the scalar and EM field equations to the first-order of the coupling constant $k_{a}$
$\nabla^{\mu}\nabla_{\mu}\phi_{(0)}-\mu_{1}^{2}\phi_{(0)}+k_{a}\left(\nabla^{\mu}\nabla_{\mu}\phi_{(1)}-\mu_{1}^{2}\phi_{(1)}-\frac{1}{2}F_{\mu\nu}^{(0)}{}^{*}F^{\mu\nu}_{(0)}\right)+\mathcal{O}\left(k_{a}^{2}\right)=0~{},$ (7)
@@
which clearly shows that at zero-order the axion cloud and the EM field obey their own field equations respectively, and at first-order the existence of a background EM field can give a substructure on the axion cloud while the existence of both axion cloud and a background EM field can induce an extra EM field in order of $k_{a}$. The first-order backreaction on the axion cloud has been studied in [27]. In this work, we focus on the EM field induced from the axion cloud in the presence of a background EM field.
@@
which describes the equilibrium state of the black hole absorbing charges and discharging processes from the surroundings of astrophysical environments such as the accretion disk, plasma and the interstellar medium [16]. It is worth noting that this extended Wald solution is still an exact solution in Kerr geometry.
For the axion cloud part, the bound states in Kerr metric can be obtained analytically in approximation when the fine structure constant $\alpha_{1}\equiv GM\mu_{1}/c=M\mu_{1}\ll 1$ in separable form as
@@
$\omega_{R}$ $\simeq$ $\mu_{1}\left(1-\frac{\alpha_{1}^{2}}{2\left(n+l+1\right)^{2}}\right)~{},$ (15)
$\omega_{I}$ $\simeq$ $\frac{1}{\gamma_{nlm}M}\left(\frac{am}{M}-2\mu_{1}r_{+}\right)\alpha_{1}^{4l+5}$ (16)
where $\gamma_{nlm}$ is a constant, $n$ is the principal quantum number, and $l$ corresponds to the orbital angular momentum quantum number. The condition for the superradiance to occur $\mu_{1}<\frac{am}{2Mr_{+}}=m\Omega_{H}$ is equivalent to $\alpha_{1}<\frac{am}{2r_{+}}=\frac{ma}{2M\left(1+\sqrt{1-a^{2}/M^{2}}\right)}$.
@@
It is necessary to clarify the difference of the EM frequencies. Turning back to the original Klein-Gordon equation (3), an axion field evolving as $\phi\sim e^{-i\omega t+im\varphi}$ would lead to EM field as $F_{\mu\nu}\sim e^{-i\omega_{EM}t+im_{EM}\varphi}$ at half frequency of the axion field $\omega_{EM}=\omega/2$. While if we look at the EM field equation (4), we cannot get a direct relation between the axion frequency and the EM frequency.
In quantum stimulated case [23], the half frequency was explained as the axion decay channel is into two photons resulting the photon energy at $E=m_{a}/2$. Also, the numerical evolution results [20, 9] gave EM oscillation frequency at half the axion mass frequency for supercritical couplings. In [27], the origin of the instability was analyzed in flat space through the Mathieu equation, which predicts instabilities for $\omega_{EM}=\mu_{1}/2,\mu_{1},3\mu_{1}/2,...$. However in [9] the frequency peaks at $\mu_{1},2\mu_{1},...$ were absent, and for high axionic couplings in dense plasmas $\omega_{p}>\mu_{1}/2$ the frequency centers at $\omega_{\rm EM}=\mu_{1}$ instead of the usual $\omega_{\rm EM}=\mu_{1}/2$, suggesting that the axion decay mechanism into two photons is forbidden and changes to process $\phi+\phi\to\gamma+\gamma$. Our understanding is that the half frequency appearing in classically numerical studies is because the main energy transfer from the axion cloud to EM field happens through the Klein-Gordon equation (3), so that even in curve spacetime we can still have a simple relation observed directly from the equation as we have analyzed previously.

### v4

@@ Abstract
It is known that the axion-photon coupling can lead to quantum stimulated emission of photons and classic exponential amplification of electromagnetic (EM) fields at half the axion mass frequency, when the axion density or the coupling constant is sufficiently large. In this work, we studied the EM photon cloud induced from an axion cloud around a Kerr black hole in the first order of the coupling constant classically. In the presence of a static EM background (such as the extended Wald solution motivated by astrophysical environments), we found that an EM photon cloud emerges, oscillating at the same frequency as the axion cloud and growing exponentially in accordance with the axion cloud when the superradiant condition for the axion field is satisfied. The evolution of the EM photon cloud with time and azimuthal angle is obtained analytically while the cross-sectional distribution is solved numerically. The induced EM field exhibits symmetries that are markedly different from those of the background EM field. Consequently, the induced photon cloud forms an unstable bound configuration that emits EM waves to spatial infinity while being replenished by the axion cloud, providing a potential observational signature of both the presence of an axion cloud and axion-photon coupling.
@@ I Introduction
Recently, there have been extensive experimental and observational efforts to search for axionic imprints. Superradiant mechanisms provide a promising avenue for probing their existence through gravitational waves [9, 10, 11], as axions can form dense clouds around spinning black holes. Although tidal perturbations from a nearby binary companion can destabilize the boson cloud and even terminate superradiance [12], the binary evolution can be significantly modulated by the backreaction with reduced termination rate [13]. The superrradiance mechanism is a well established process for extracting energy from a rotating black hole [15, 14]. Penrose and Floyd [16] first demonstrated that a particle entering the ergoregion of a rotating black hole can split such that one fragment falls in with negative energy, allowing the other to escape with excess energy, thereby extracting the black hole’s rotational energy. Subsequently, Misner [17] showed that waves with frequencies $\omega<m\Omega_{H}$ can also extract rotational energy through superradiant scattering, in which an incoming wave is amplified as it scatters off a rotating black hole. Later, Press and Teukolsky [18] introduced the concept of the black hole bomb, and Teukolsky further generalized the mechanism to EM and gravitational waves [19]. Nevertheless, the Kerr solution remains stable against massless scalar, EM and gravitational perturbations without reflection.
It is well known that an isolated black hole cannot sustain an intrinsic EM field unless it carries a net electric charge. Thus, for a black hole to exhibit EM effects, some charging mechanism is required. In realistic astrophysical environments, surrounding plasma and accretion flows in the presence of large-scale magnetic fields can give rise to effective charge separation or induce a nonvanishing net charge on the black hole. Possible origins of such EM configurations include charge separation induced by magnetospheric processes such as the Blandford–Znajek mechanism, effects associated with plasma inflow, disk–magnetosphere interactions, as well as pair-production–supported magnetospheres. In the present work, however, we do not model these mechanisms explicitly and instead adopt the extended Wald solution [20] as an effective, stationary background description. Neglecting backreaction and making use of the two Killing vectors of the Kerr spacetime, Wald constructed an exact test solution of Maxwell’s equations describing a homogeneous magnetic field aligned with the black hole spin axis, which could lead to the accretion of charges from the surrounding plasma in the accretion disk or the interstellar medium. As a result, the black hole would acquire a charge and be described by a Kerr-Newman spacetime. However, at equilibrium with the original magnetic field, this EM configuration is also a source-free test solution on the Kerr metric, so that we can consider it as the EM background solution in our work. Besides, extraction of energy and charge from a black hole was studied in [21]. The superradiance effects and the multimessenger signals of an EM field were studied in [22, 11]. Recently, new polarized images from the Event Horizon Telescope (EHT) collaboration revealed significant variations in brightness and polarization patterns across 2017, 2018, and 2021, highlighting the unexpectedly evolving magnetic field structures around the black hole M87* [23].
Although the axion-photon coupling is expected to be very weak, it was suggested that the quantum stimulated emission of photons can be generated at large enough axion number [24, 25], and conjectured that blasts of light could be emitted from black hole systems [26, 27]. Recently at classical level, it was shown that EM fields can be exponentially amplified in Kerr-axion cloud system, when the amplitude of the axion cloud multiplied by the coupling constant is larger than a critical value [28]. Thereinto, a fraction of the cloud energy is transferred to the EM blasts, while the depleted axion cloud will be replenished via superradiance, leading to the spin-down of the black hole as it was found in [29]. In [30], it was also found that the presence of an axion field can cause the reversal of rotation in Kerr black holes. Even in the presence of plasma, the EM instability is still controlled by the axionic coupling [31]. Here the plasma effects are characterized by the plasma frequency $\omega_{p}$, which acts as an effective mass scale for EM waves and governs plasma screening. Interestingly, a new class of superradiant instabilities of axion bound states in neutron star magnetospheres has been identified [32], in which axion-photon mixing allows the extraction of rotational energy from the neutron star, since rotation and a dissipation channel are enough to trigger superradiance [15].
In a flat-spacetime analysis with a uniform oscillating axion background, the instability appears through the Mathieu equation, originating from the homogeneous part of the equation [33]. In contrast, the flat-spacetime analysis in a homogeneous EM background showed that the axion-photon coupling modifies the dispersion relation, and that a sufficiently strong electric field can trigger an instability with a growing mode. This latter instability arises from the inhomogeneous (sourced) part of the EM field equation. In fact, in the presence of external EM fields, photons can mix with low-mass bosons if they couple via a two-photon vertex, analogous to neutrino-flavor-mixing effects [34]. In addition, both axion-photon mixing around charged black holes and dark photon-photon mixing around neutral black holes were studied in the presence of plasma [35], in which the flat spacetime analysis in the presence of a magnetic field gave the mixing matrixes with non-diagonal terms. Moreover, superradiant instabilities by photons confined around black holes in the presence of astrophysical plasmas were investigated under the plasma-photon interaction, where quasi-bound EM states were found to be superradiantly unstable for rotating black holes though it probably be quenched due to plasma blowout in the nonlinear evolution [36, 37, 38]. From the observational side, with the polarimetric measurements of M87* by the EHT collaboration stringent constraints on the axion-photon coupling constant in the mass range of $\sim\left(10^{-21}-10^{-20}\right)$eV were given, based on the azimuthal distribution of the electric vector position angle [39]. Furthermore, the deviation of photon geodesics due to oscillating metric perturbations induced from the superradiant clouds was predicted in [40].
@@
In this work we studied the EM photon cloud induced from an axion cloud via the axion-photon coupling surrounding a Kerr black hole. In the first order of the coupling constant, we found that a background EM field is necessary to have a first-order induced EM structure, and if it is stationary then the frequency of the induced photon cloud will be the same as the frequency of the axion cloud. It means that it grows or decays in accordance with the axion cloud when the superradiant condition for the axion field is satisfied or not, while the oscillating frequency is also the same. It is well known that the timescale of the superradiant instability is much larger than the timescale of the oscillation, therefore we can ignore the growth of the amplitude and focus on the oscillating part.
We found that no threshold on the coupling constant or on the amplitude of the axion cloud is required for the emergence of this oscillating and growing photon cloud. The instability is inherited from the axion cloud in the presence of an EM background. This is in contrast to the case without a background EM field, where a threshold for the amplitude of the axion cloud multiplied by the coupling constant and an initial EM pulse are both required; otherwise, no instability develops and the initial EM fluctuation decays exponentially. While for the dark photon-photon coupling case, a dark photon cloud can transfer its energy to form a rotating EM field directly without any external EM fields [22], since no background EM fields are necessary for the mixing to occur [35].
@@
The work is organized as follows. Firstly in Sec. II we study the basic field equations in the first-order of the coupling constant, then in Sec. III we show how an EM photon cloud is induced from the axion cloud in the presence of an EM background and discuss symmetries. Moreover, in Sec. IV we study the induced electric and magnetic fields respectively. Finally in Sec. V we conclude.
@@ II First-order Analysis of the Axion-Photon Coupling
$\mathcal{L}_{m}=-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}-\frac{1}{2}\nabla_{\mu}\phi\nabla^{\mu}\phi-\frac{1}{2}\mu^{2}\phi^{2}-\frac{k_{a}}{2}\phi F_{\mu\nu}{}^{*}F^{\mu\nu}~,$ (2)
where the axion field $\phi$ is a real pseudo scalar field with mass $m_{a}\equiv\mu\hbar$, $k_{a}$ is the coupling constant, $F_{\mu\nu}\equiv\nabla_{\mu}A_{\nu}-\nabla_{\nu}A_{\mu}$ is the electromagnetic field strength tensor and $^{*}F^{\mu\nu}\equiv\frac{1}{2}\epsilon^{\mu\nu\rho\sigma}F_{\rho\sigma}$ is the dual field strength tensor. Here $\epsilon^{\mu\nu\rho\sigma}\equiv\frac{1}{\sqrt{-g}}E^{\mu\nu\rho\sigma}$ and $E^{\mu\nu\rho\sigma}$ denotes the totally antisymmetric Levi-Civita symbol with $E^{0123}=+1$. We use geometrized units $c=G=1$ unless otherwise specified.
@@
$\nabla^{\mu}\nabla_{\mu}\phi-\mu^{2}\phi-\frac{k_{a}}{2}F_{\mu\nu}{}^{*}F^{\mu\nu}=0~,$ (3)
@@
In our study, the Kerr geometry is fixed as the background, since the backreactions of both axion cloud and the EMs on the metric are in the second-order of the fluctuations. To clarify the hierarchy, we introduce small quantities $\eta$ and $e$ to represent the magnitudes of the axion cloud and the EM field respectively as $F_{\mu\nu}\to eF_{\mu\nu}$ and $\phi\to\eta\phi$, and the energy-momentum tensor becomes
$T_{\mu\nu}$ $\to$ $-\frac{1}{2}g_{\mu\nu}\left(\frac{1}{2}e^{2}F_{\alpha\beta}F^{\alpha\beta}+\eta^{2}\nabla_{\sigma}\phi\nabla^{\sigma}\phi+\mu^{2}\eta^{2}\phi^{2}+k_{a}e^{2}\eta\phi F_{\alpha\beta}{}^{*}F^{\alpha\beta}\right)$~ (7)
$+e^{2}F^{\sigma}{}_{\mu}F_{\sigma\nu}+\eta^{2}\nabla_{\mu}\phi\nabla_{\nu}\phi+4k_{a}e^{2}\eta\phi F^{\sigma}{}_{\nu}{}^{*}F_{\sigma\mu}~,$
showing that the magnitudes of the backreaction on the metric are of the orders $\sim e^{2},~\eta^{2},~\mu^{2}\eta^{2},~k_{a}\eta e^{2}$. The other field equations become
$\nabla^{\mu}\nabla_{\mu}\phi-\mu^{2}\phi-\frac{k_{a}e^{2}}{2\eta}F_{\mu\nu}{}^{*}F^{\mu\nu}=0~,$ (8)
$\nabla_{\nu}F^{\mu\nu}+2k_{a}\eta\nabla_{\nu}\phi{}^{*}F^{\mu\nu}=0~,$ (9)
which yields the conditions $\mu^{2}\eta\gg k_{a}e^{2}$ and $k_{a}\eta\ll 1$ when we consider the axion cloud and the EM field as background fields, with the effects of their interaction appearing only at higher orders.
Using the typical field amplitude in the boson cloud $\langle\phi^{2}\rangle\sim 8\tilde{a}\alpha^{5}M_{\rm Pl}^{2}$ [41], and the limit for the largest magnetic field $B\sim 4\times 10^{18}\left(M/M_{\odot}\right)^{-1/2}{\rm Gauss}$ that can be supported by an accretion disk around astrophysical BHs [42], we obtain the first condition
$k_{a}\ll\frac{160.705\sqrt{\tilde{a}}\alpha^{9/2}}{M_{\rm BH}/M_{\odot}}~{\rm GeV}^{-1}$~ (11)
that the background solution of the axion cloud without EM field is valid, and the relevant axion mass is around
$m_{a}\sim\frac{\alpha}{M_{\rm BH}/M_{\odot}}1.3\times 10^{10}~{\rm eV}~.$ (12)
Here we adopt natural units, setting $\hbar=c=1$ and $4\pi\epsilon_{0}=1$, in order to facilitate comparison with other constraints on the coupling constant. For instance, in the case of the supermassive black hole M87*, taking $\tilde{a}=0.8$ and $\alpha=0.2$, we have $k_{a}\ll 1.58233\times 10^{-11}~{\rm GeV}^{-1}$. In contrast, for a primordial black hole with $M/M_{\odot}=0.2$, $\tilde{a}=0.5$ and $\alpha=0.2$, the condition becomes much less restrictive, $k_{a}\ll 0.406555~{\rm GeV}^{-1}$. We will verify the second validity condition of the perturbative framework for the EM field equation, after we obtain the amplitude of the first-order EM field.
We now return to the original field equations and expand them to first order in the coupling constant $k_{a}$
$\nabla^{\mu}\nabla_{\mu}\phi_{(0)}-\mu^{2}\phi_{(0)}+k_{a}\left(\nabla^{\mu}\nabla_{\mu}\phi_{(1)}-\mu^{2}\phi_{(1)}-\frac{1}{2}F_{\mu\nu}^{(0)}{}^{*}F^{\mu\nu}_{(0)}\right)+\mathcal{O}\left(k_{a}^{2}\right)=0~,$ (13)
@@
which clearly shows that to zeroth order the axion cloud and the EM field obey their own field equations respectively, while at first order the background EM field gives a substructure on the axion cloud and the coexistence of the axion cloud and the background EM field generates an extra EM field of order $k_{a}$. The first-order backreaction on the axion cloud has been studied in [33]. In this work, we focus on the EM field induced from the axion cloud in the presence of a background EM field.
@@
which describes the equilibrium state of the black hole absorbing charges and discharging processes from the surroundings of astrophysical environments such as the accretion disk, plasma and the interstellar medium [20]. It is worth noting that this extended Wald solution is still a source-free test solution on the Kerr geometry.
For the axion cloud part, the bound states in Kerr metric can be obtained analytically in approximation when the fine structure constant $\alpha\equiv GM\mu/c=M\mu\ll 1$ in separable form as
@@
$\omega_{R}$ $\simeq$ $\mu\left(1-\frac{\alpha^{2}}{2\left(n+l+1\right)^{2}}\right)~,$ (21)
$\omega_{I}$ $\simeq$ $\frac{1}{\gamma_{nlm}M}\left(\frac{am}{M}-2\mu r_{+}\right)\alpha^{4l+5}$ (22)
where $\gamma_{nlm}$ is a constant, $n$ is the principal quantum number, and $l$ corresponds to the orbital angular momentum quantum number. The condition for the superradiance to occur $\mu<\frac{am}{2Mr_{+}}=m\Omega_{H}$ is equivalent to $\alpha<\frac{am}{2r_{+}}=\frac{ma}{2M\left(1+\sqrt{1-a^{2}/M^{2}}\right)}$.
@@
It is necessary to clarify the difference in the EM frequencies. Returning to the original Klein-Gordon equation (3), an axion field evolving as $\phi\sim e^{-i\omega t+im\varphi}$ would lead to the EM field as $F_{\mu\nu}\sim e^{-i\omega_{\rm EM}t+im_{\rm EM}\varphi}$ at half frequency of the axion field $\omega_{\rm EM}=\omega/2$. However, if we look at the EM field equation (4), a direct relation between the axion and the EM frequencies cannot be obtained.
In the quantum stimulated case [24], the half frequency was explained by the axion decay channel $\phi\to 2\gamma$, where each photon carries energy $E=m_{a}/2$. Numerical simulations [28, 31] also showed that the EM oscillation frequency appears at half the axion mass frequency for supercritical couplings. In [33], the origin of the instability was analyzed in flat spacetime through the Mathieu equation, which predicts instabilities for $\omega_{\rm EM}=\mu/2,\mu,3\mu/2,$.... Whereas in [31] the frequency peaks at $\mu,2\mu,$... were absent, and for high axionic couplings in dense plasmas with plasma frequency $\omega_{p}>\mu/2$, the EM spectrum instead centers at $\omega_{\rm EM}=\mu$ rather than at the usual $\omega_{\rm EM}=\mu/2$. This behavior suggests that the standard axion decay channel into two photons is kinematically forbidden and changes to process $\phi+\phi\to\gamma+\gamma$.